OpenStudy (anonymous):

A bullet of mass 20g is fired into a 3kg wooden block hung on a rope. the block, with bullet embedded, swings up so that it makes a 22degree arc before falling back down. what speed was the bullet fired? (it's a ballistic pendulum. 1meter is the string

OpenStudy (anonymous):

 It's a two part problem. First conservation of evergy, since you know how high the block went, and you know it initially had only kinetic energy, and when it stops at the top of the arc has only potential energy \[E_i=E_f\] \[\frac{1}{2}(M+m)v^2 = (M+m)gh\] solve for v Then you use that v in a conservation of momentum problem, knowing that the block was initially at rest, and that after the collision the block and bullet are stuck together \[p_i=p_f\] \[mv_i = (M+m)v\] and solve for vi (of the bullet)

OpenStudy (anonymous):

where M is the mass of the block and m is the mass of the bullet.

OpenStudy (anonymous):

oh okay thank you

OpenStudy (anonymous):

Does that all make sense? ^_^

OpenStudy (anonymous):

what is the unit of this?

OpenStudy (anonymous):